# Diffusion approximation of the stationary distribution of a two-level single server queue

### v6

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic. This limit corresponds to its diffusion approximation. It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters. This result is proved using the so called BAR approach studied in [2, 3, 10, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
Keywords: single server queue; level dependent arrival and service; stationary distribution; heavy traffic; diffusion approximation; Palm distribution; stochastic integral equation; multi-level SRBM.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured. Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
There is another motivation of this study. In general, theoretical study of diffusion approximation is hard for state-dependent queues. For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see Mandelbaum and Pats [8], Yamada [14]), while the arrival process and service rate discontinuously change in the present model. Thus, theoretical study for its diffusion approximation is challenging.
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see Kingman [7]). Furthermore, Miyazawa [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see Harrison [5]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [2, 3, 10, 12]. This verification, Theorem 3.1, requires either some limitation on system parameters or for the workload brought by a customer to be exponentially distributed. The proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory.
@@
This paper is made up by five sections. We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1. A main result, Theorem 3.1, and the exponentially distributed case, Proposition 3.1, are presented in Section 3.2. In Section 4, we prove Theorem 3.1. Finally, in Section 5, we intuitively discuss two topics. In Section 5.1, we consider a process-limit for the 2-level $GI/G/1$ queue in heavy traffic. In Section 5.2, we intuitively verify the conjecture in Remark 3.1.
@@ 2 Framework for the 2-level GI/G/1 queue
According to Miyazawa [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
@@ 3.1 Assumptions
Under the stationary assumption, $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$. Similarly, $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[(c^{(r)}_{1}\wedge c^{(r)}_{1})T_{s}]<\infty$. ∎
@@ 3.2 Main results
$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.1)
@@
$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}>\ell^{(r)}_{0}),$
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}>\ell^{(r)}_{0}).$
@@
    For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
@@
    If either $\lambda_{1}=\lambda_{2}$ (equivalently, $c_{1}=c_{2}$) or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then $\alpha^{(r)}_{e}$ converges to $\widetilde{\alpha}_{e}$ and $\nu^{(r)}$ weakly converges to $\widetilde{\nu}$, namely
    $\lim_{r\downarrow 0}\nu^{(r)}(B)=\widetilde{\nu}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$ (3.11)
@@
Under either one of the assumptions in (ii) of Theorem 3.1,
@@
Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue.
@@
Under the settings of Theorem 3.1, the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1. This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of more than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). This Markov chain will be detailed in Appendix A.2.
4 Proofs
In this section we prove Theorem 3.1. A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [2, 3, 10]. Recently, this approach is further studied in [12].
The proof of Theorem 3.1 is composed of three parts. We first introduce the BAR approach. We then prepare auxiliary results, and finally prove Theorem 3.1.
@@ 4.1 Palm distribution and BAR
${\mathcal{H}}f({\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f({\boldsymbol{x}})-(c_{1}1(1\leq x_{1}\leq\ell_{0})+c_{2}1(x_{1}>\ell_{0}))\frac{\partial^{+}}{\partial x_{3}}f({\boldsymbol{x}}),\qquad{\boldsymbol{x}}\in S,$
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
@@
The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short.
@@
    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distribution $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. So, $X^{(r)}(0-)$ can be considered as a random variable generated by $\mathbb{P}^{(r)}_{e}$. However, one may be uncomfortable for this interpretation. In this case, please consider the process $X(t)$ to start at time $-1$ subject to the stationary distribution, then the distribution of $X(\cdot)=\{X(t);t\geq 0\}$ is unchanged because $X(\cdot)$ is a stationary process. In either way, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions.
@@
For each integer $k\geq 0$, let $f({\boldsymbol{x}})=x_{1}\wedge(k+1)$ for ${\boldsymbol{x}}\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
@@
$\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
@@
$\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{E}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$. Because of this equality, the above formula also yields (4.3). We next let $f({\boldsymbol{x}})=x_{2}$, and apply (4.1), then
@@
because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.4). Similarly, let $f({\boldsymbol{x}})=x_{3}$, then
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 4.3.
4.2 Preliminary lemmas

### v7

@@ Abstract
We consider a single server queue which has a threshold to change its arrival process and service rate by its queue length, which is referred to as a two-level single server queue. This model is motivated by an energy saving problem for a single server queue whose arrival process and service rate may be changed. To get its performance in tractable form, we study the limit of the stationary distribution of the queue length in this two-level queue under scaling in heavy traffic. This limit corresponds to its diffusion approximation. It is shown that this limiting distribution is truncated exponential (or uniform if the drift is null) below the threshold level and exponential above it under suitably chosen system parameters and generally distributed inter-arrival times and workloads brought by customers. This result is proved under a mild limitation on arrival parameters using the so called BAR approach studied in [3, 4, 10, 12]. We also intuitively discuss about a diffusion process corresponding to the limit of the stationary distribution under scaling.
@@ 1 Introduction
So far, we are interested in the performance of this 2-level $GI/G/1$ queue over long time, for example the mean queue length. However, it is hard to derive such a characteristic in an analytically tractable form in general, while, for the special case that all the distributions are exponential, its analytic expression is available (see Appendix A.2), but the variability of the arrival times and service amounts are not well captured in this exponential case. Because of these facts, an approximation by diffusion is helpful, so we study it through the limit of the stationary distributions of the scaled queue lengths of this 2-level $GI/G/1$ queues in heavy traffic, where the scale factor is chosen so that the limit is well obtained as a probability distribution. This limiting distribution corresponds to the stationary distribution of a one-dimensional reflected Brownian motion whose drift and variance are discontinuously changed at the level corresponding to the threshold, which is recently studied in [11, 13]. Thus, we will indirectly consider a diffusion approximation for the 2-level $GI/G/1$ queue.
There is another motivation of this study. In general, theoretical study of diffusion approximation is hard for state-dependent queues. For example, there are some theoretical studies about them, but they assume that the arrival and service processes are of exponential types, and their rates are smoothly changed (e.g., see [8, 15]), while the arrival process and service rate discontinuously change in the present model. Thus, theoretical study for its diffusion approximation is challenging.
As is well known, the scaling limit of the stationary distribution is exponential for the standard $GI/G/1$ queue in heavy traffic (e.g., see [7]). Furthermore, [12] shows that the limiting distribution is truncated exponential (or uniform for the null drift case) for the $GI/G/1$ queue with a finite waiting room. Both distributions agree with the stationary distributions of the corresponding reflected Brownian motions (see [6]). Thus, one may conjecture that the limiting distribution for the 2-level $GI/G/1$ queue is truncated exponential (or uniform for the null drift case) below the threshold level and exponential above it. This conjecture can be directly verified if all the distributions are exponential (see Proposition 3.1), but this is exceptional. Under the generally distributed assumptions, we verify it, employing the BAR approach studied in [3, 4, 10, 12]. This verification, Theorem 3.1, requires a mild limitation on the parameters of the arrival processes (see (3.f)). Its proof needs some knowledge about Palm distributions and a truncation technique in computations, but does not require any deeper theory. However, the proof has a complicated structure, so we split it into many components so that they are orthogonal as long as possible. This is the reason why we spend two sections for the proof.
@@
This paper is made up by six sections. We first introduce a general framework for a Markov process describing the 2-level $GI/G/1$ queue in Section 2, then provide notations and assumptions for a sequence of the 2-level $GI/G/1$ queues in heavy traffic in Section 3.1. A main result, Theorem 3.1, and the exponentially distributed case, Proposition 3.1, are presented in Section 3.2. In Section 4, we introduce the so called BAR approach, and apply it to the 2-level $GI/G/1$ queue. Using these results, Theorem 3.1 is proved in Section 5. Finally, in Section 6, we intuitively discuss two topics. In Section 6.1, we consider a process-limit for the 2-level $GI/G/1$ queue in heavy traffic. In Section 6.2, we intuitively verify the conjecture in Remark 3.2.
@@ 2 Framework for the 2-level $GI/G/1$ queue
According to [12], we take stochastic basis $(\Omega,{\mathcal{F}},\mathbb{F},\mathbb{P})$ with time-shift operator semi-group $\Theta_{\bullet}\equiv\{\Theta_{t};t\geq 0\}$ to jointly consider the Markov process $X(\cdot)$ and the counting processes $N_{e}(\cdot)$ and $N_{s}(\cdot)$, where $\mathbb{F}\equiv\{{\mathcal{F}}_{t};t\geq 0\}$ is a filtration generated by $X(\cdot)$, that is, ${\mathcal{F}}_{t}=\sigma(\{X(u);0\leq u\leq t\})$, and $\{\Theta_{t};t\geq 0\}$ is a operator semi-group (namely, $\Theta_{s+t}=\Theta_{s}\circ\Theta_{t}$ for $s,t\geq 0$) on $\Omega$ satisfying the following two conditions.
@@ 3.1 Assumptions
Most of our arguments are valid under the assumptions (3.a)–(3.e), but we require the following condition for their final step to determine the limiting distribution of the scaled queue length in heavy traffic.
  • (3.f)
    $\sigma_{e,1}^{2}=\sigma_{e,2}^{2}$ if $\lambda_{1}=\lambda_{2}$.
Remark 3.1.
The condition (3.f) comes out from technical requirements for the proof of Theorem 3.1. It is a mild limitation because no condition is needed for $\lambda_{1}\not=\lambda_{2}$, but we think that it is likely not needed. However, removing (3.f) seems to be beyond the framework of this paper. See Remark 5.1 for its details.
@@
Assume the assumptions (3.a)–(3.e) and the stationarity of $X^{(r)}(\cdot)$, then $\alpha^{(r)}_{e}$ and $\alpha^{(r)}_{s}$ are finite.
@@
Let $N^{(r)}_{e,0}(\cdot)$ be the renewal process whose inter-counting times are $T_{e,1}(n)\wedge T_{e,2}(n)$ for $n\geq 1$. Obviously, $\mathbb{E}[N^{(r)}_{e}(t)]\leq\mathbb{E}[N^{(r)}_{e,0}(t)]$ under the stationary framework. Hence, $\alpha^{(r)}_{e}=\mathbb{E}[N^{(r)}_{e}(1)]\leq\mathbb{E}[N^{(r)}_{e,0}(1)]=1/\mathbb{E}[T_{e,1}\wedge T_{e,2}]<\infty$. Similarly, let $N^{(r)}_{s,0}(\cdot)$ be the renewal process whose inter-counting times are $((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1}))^{-1}T_{s}(n)$ for $n\geq 1$, then $\alpha^{(r)}_{s}=\mathbb{E}[N^{(r)}_{s}(1)]\leq\mathbb{E}[N^{(r)}_{s,0}(1)]=1/\mathbb{E}[((c^{(r)}_{1})^{-1}\wedge(c^{(r)}_{1})^{-1})T_{s}]<\infty$. ∎
@@ 3.2 Main results
$\nu^{(r)}(B)=\mathbb{P}(rL^{(r)}\in B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}).$
@@
$\nu^{(r)}_{1}(B)=\mathbb{P}(rL^{(r)}\in B;L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2}(B)=\mathbb{P}(rL^{(r)}\in B;\ell^{(r)}_{0}<L^{(r)}),$
$\nu^{(r)}_{1|cd}(B)=\mathbb{P}(rL^{(r)}\in B|L^{(r)}\leq\ell^{(r)}_{0}),\quad\nu^{(r)}_{2|cd}(B)=\mathbb{P}(rL^{(r)}\in B|\ell^{(r)}_{0}<L^{(r)}).$
@@
    For each $i=1,2$, if $\nu^{(r)}_{i}(\mathbb{R}_{+})$ is bounded away from 00 as $r\downarrow 0$, that is,
    $\liminf_{r\downarrow 0}\varphi^{(r)}_{i}(0)>0,\qquad i=1,2,$ (3.6)
    then $\nu^{(r)}_{i|cd}$ weakly converges to $\widetilde{\nu}_{i|cd}$, namely,
@@
    If either $\lambda^{(r)}_{1}-\lambda^{(r)}_{2}=O(r)$, equivalently, $c^{(r)}_{1}-c^{(r)}_{2}=O(r)$, or $T^{(r)}_{s}$ is exponentially distributed for all $r\in(0,1]$, then
    $\lim_{r\downarrow 0}\nu^{(r)}(B)=\widetilde{\nu}(B),\qquad B\in{\mathcal{B}}(\mathbb{R}_{+}),$ (3.11)
    $\lim_{r\downarrow 0}\alpha^{(r)}_{e}=\widetilde{\alpha}_{e}.$ (3.12)
@@
Under the assumptions (3.a)–(3.f) and either one of the assumptions in (ii) of Theorem 3.1,
@@
Let us consider Theorem 3.1 when $T_{e,1}^{(r)},T_{e,2}^{(r)}$ and $T_{s}^{(r)}$ are all exponentially distributed. We call this model the $r$-th 2-level $M/M/1$ queue. Note that (3.f) is automatically satisfied in this exponential model.
@@
Under the assumptions (3.a)–(3.e), the scaled stationary distribution $\nu^{(r)}_{\exp}$ of the $r$-th 2-level $M/M/1$ queue weakly converges to $\widetilde{\nu}_{\exp}$ as $r\downarrow 0$, where $\widetilde{\nu}_{\exp}$ is defined as $\widetilde{\nu}_{\exp}(B)=\widetilde{\nu}_{1|cd}(B)\widetilde{A}_{\exp,1}+\widetilde{\nu}_{2|cd}(B)\widetilde{A}_{\exp,2}$ for $B\in{\mathcal{B}}(\mathbb{R}_{+})$ in which
@@
Of course, Proposition 3.1 is the special case of (ii) of Theorem 3.1, so no proof is needed. However, for a double check, we give its direct proof in Appendix A.2, which is helpful to convince (ii) of Theorem 3.1. This proof is challenging because the exponential model requires a Markov chain with two-dimensional state space to correctly describe the case that the new arrival will find a system with less than $\ell_{0}$ after multiple consecutive departures when the previous arrival found a queue length of greater than $\ell_{0}$ (see the arrival at time $t^{(r)}_{e}(8)$ in Figure 3.1). This Markov chain will be detailed in Appendix A.2.
4 BAR’s for the 2-level $GI/G/1$ queue
In this section, we prepare auxiliary notations and results for proving Theorem 3.1. A major difficulty comes from the fact that $L^{(r)}(\cdot)$ is not a Markov chain and we have to work with $X^{(r)}(\cdot)$. Instead of directly computing the stationary distribution, we study it through a stationary equation, called a basic adjoint relationship, BAR in short, using Palm distributions, which will be defined below. This method is called a BAR approach, and has been used in [3, 4, 10]. Recently, this approach is further studied in [12].
Throughout this section, we assume (3.a)–(3.e), but (3.f) is not assumed. The proof of Theorem 3.1 will be done in Section 5.
@@ 4.1 Palm distribution and BAR
${\mathcal{H}}f(z,{\boldsymbol{x}})=-\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}})-(c_{1}1(1\leq z\leq\ell_{0})+c_{2}1(z>\ell_{0}))\frac{\partial^{+}}{\partial x_{2}}f(z,{\boldsymbol{x}}),\quad(z,{\boldsymbol{x}})\in S,$
where ${\boldsymbol{x}}=(x_{1},x_{2})\in\mathbb{R}_{+}^{2}$.
@@
The formula (4.1) is a special case of the rate conservation law (e.g., see [9]), and referred as a basic adjoint relationship, BAR in short, then we call analysis based on this BAR as a BAR approach. A major difficulty in this BAR approach is to evaluate the expectations under the Palm distributions (the last two terms in (4.1)). This evaluation will be done for the 2-level $GI/G/1$ queue in Section 4.3.
@@
    $X^{(r)}(0-)$ is not defined under $\mathbb{P}$, but its distribution is well defined under the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{e}$ because the definition (4.1) only uses $X^{(r)}(t)$ for $t>0$ under $\mathbb{P}$. So, $X^{(r)}(0-)$ is a random variable well defined under $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$. Hence, $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ in (4.1) are well defined under the Palm distributions. If we define the stationary process $X^{(r)}(\cdot)$ on the time axis (,+)($\infty$,$\infty$)( - ∞ , + ∞ ), $\Delta_{e}f(X^{(r)}(0))$ and $\Delta_{s}f(X^{(r)}(0))$ are well defined even under $\mathbb{P}$.
@@
For each integer $k\geq 0$, let $f(z,{\boldsymbol{x}})=z\wedge(k+1)$ for $(z,{\boldsymbol{x}})\in S$, then obviously $f\in C_{b}^{p}(S)$. Since $\Delta_{e}f(X^{(r)}(u)=1(L^{(r)}(u-)\leq k,N^{(r)}_{e}(u)-N^{(r)}_{e}(u-)=1)$ and $\Delta_{s}f(X^{(r)}(u)=-1(L^{(r)}(u)\leq k,N^{(r)}_{s}(u)-N^{(r)}_{s}(u-)=1)$, we have
@@
$\quad=\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right].$
@@
$\alpha^{(r)}_{e}\mathbb{P}^{(r)}_{e}\left[L^{(r)}(0-)\leq k\right]-\alpha^{(r)}_{s}\mathbb{P}^{(r)}_{s}\left[L^{(r)}(0)\leq k\right]=0,\qquad k\geq 0.$
Hence, letting $k\to\infty$, we have $\alpha^{(r)}_{e}=\alpha^{(r)}_{s}$. Because of this equality, the above formula also yields (4.3). We next let $f(z,{\boldsymbol{x}})=x_{1}$, and apply (4.1), then
@@
because $(R^{(r)}_{e})^{\prime}(t)=-1$. This proves (4.4). Similarly, let $f(z,{\boldsymbol{x}})=x_{2}$, then
@@
To prove Theorem 3.1, we will fully use the BAR (4.1) replacing $\alpha^{(r)}_{s}$ by $\alpha^{(r)}_{e}$, which is possible by Lemma 4.1. We first apply an exponential type of a test function to this BAR, and derive various properties on it under the stationary distribution $\mathbb{P}$ and the Palm distributions $\mathbb{P}^{(r)}_{e}$ and $\mathbb{P}^{(r)}_{s}$ in Section 4.2, then Theorem 3.1 is proved in Section 5.
4.2 Asymptotic BAR for the pre-limit process $X^{(r)}(\cdot)$